What is the ratio of those who prefer black coffee to those who prefer nonblack coffee.Among 100 Coffee drinkers, 36 said that they preferred their coffee black and the rest did not prefer their coffee black. : What is the ratio of those who prefer black coffee to those who prefer nonblack coffee.Among 100 Coffee drinkers, 36 said that they preferred their coffee black and the rest did not prefer their coffee black.
Answer by orca(336)

The number of people who prefer black coffee is 36.
The number of people who prefer nonblack coffee is 100 - 36 = 64.
Their ratio is 36 : 64 = 4*9/4*16 = 9/16 = 9 : 16.